If r varies jointly as s and t, when r=90, s=8 and t=15. find r when s=7 and t=12.
Answer by Alan3354(69443)

If r varies jointly as s and t, when r=90, s=8 and t=15. find r when s=7 and t=12.
----------
You didn't specify, so I'll assume r varies directly, not inversely.
r = kst (k is some number)
Use the data given to find k
90 = k*8*15 = 120k
k = 90/120 = 3/4
-------------------
@ s=7, t=12
r = (3/4)*7*12
r = 63